Determine the specific heat capacities at constant volume and constant pressure for methane, given that its molecular weight is 16 and the adiabatic index of expansion and compression is 2.3 the universal gas is 8314 j Kmol-1/K-1
1
Expert's answer
2018-06-28T09:12:08-0400
ɤ = Cp/Cv = 2.3
Cp = 2.3Cv
Cv = Cp – nR
Cv = 2.3Cv – nR
R = 1.3Cv
Cv = 0.77R
Cp = 2.3*0.77R = 1.77R